OpenStudy (anonymous):

Solve: 3(x + 4) = 5(x − 2)

mathslover (mathslover):

remember the distributive property : \[\huge{a(b+c)=ab+ac}\] \[\huge{3(x+4)=5(x-2)}\] \[\huge{3x+3(4)=5(x)-5(2)}\] \[\huge{3x+12=5x-10}\] \[\huge{12+10=5x-3x}\] can u solve this further ?

OpenStudy (anonymous):

you solve for x

OpenStudy (anonymous):

3(x+4)=5(x-2) 3x+12=5x-10 5x-3x=12+10 2x=2 x=1

OpenStudy (anonymous):

3x + 12 = 5x - 10 3x -5x = -10-12 -2x=-22 x=11 :)